OpenStudy (anonymous):

^4sqrt(v^36) it says v^8 is incorrect

OpenStudy (anonymous):

V^9

OpenStudy (anonymous):

duh, im such a ditz

OpenStudy (anonymous):

oh wait it says thats not right either

OpenStudy (anonymous):

What is the question exactly?

OpenStudy (anonymous):

\[\sqrt[4]{v^{36}}=v^\frac{36}{4}=v^9\]